### v1

# Mock Eisenstein series associated to partition ranks

Abstract.
In this paper, we introduce a new class of mock Eisenstein series, describe their modular properties, and write the partition rank generating function in terms of so-called partition traces of these. Moreover, we show the Fourier coefficients of the mock Eisenstein series are integral and we obtain a holomorphic anomaly equation for their completions.
Key words and phrases:
completions, holomorphic anomaly equations, (mock) Jacobi forms, partition traces, ranks.
2020 Mathematics Subject Classification:
11F03, 11F11, 11F37, 11F50, 11P82
1. Introduction and statement of results
A partition of $n\in\mathbb{N}_{0}$ is a weakly decreasing sequence of positive integers that sum to $n$. We denote by $p(n)$ the number of partitions of $n$. Recall the famous Ramanujan congruences
[…]
To explain the first two, Dyson [10] introduced the rank of a partition $\lambda$, which is defined as
$\mathrm{rank}(\lambda):=\text{largest part of }\lambda-\text{number of parts of }\lambda.$
Dyson conjectured that reducing the rank (mod $5$) (resp. $7$) divides the partitions of $45n+$ 4 (resp. $57n+$ 5) into $5$ (resp. $7$) sets of equal size. This conjecture was proven by Atkin and Swinnerton-Dyer [5]. In the same paper, Dyson also conjectured the existence of another statistic, which he called the “crank” and which should explain all three partition congruences. Garvan [12] found a crank for vector partitions and Andrews–Garvan [4] defined a crank for ordinary partitions. Letting $o(\lambda)$ denote the number of ones in a partition $\lambda$, and $\mu(\lambda)$ the number of parts strictly larger than $o(\lambda)$, the crank is defined as
$\mathrm{crank}(\lambda):=\begin{cases}\text{largest part of }\lambda&\text{if }o(\lambda)=0,\\ \mu(\lambda)-o(\lambda)&\text{if }o(\lambda)>0.\end{cases}$
Let $N(m,n)$ denote the number of partitions of $n$ with rank $m$. Its generating function is
$R(\zeta;q):=\sum_{\begin{subarray}{c}n\geq 0\\ m\in\mathbb{Z}\end{subarray}}N(m,n)\zeta^{m}q^{n}=\sum_{n\geq 0}\frac{q^{n^{2}}}{(\zeta q)_{n}\left(\zeta^{-1}q\right)_{n}},$ (1.1)
(see [5]) where $(a)_{n}=(a;q)_{n}:=\prod_{j=0}^{n-1}(1-aq^{j})$ for $n\in\mathbb{N}_{0}\cup\{\infty\}$. Let $M(m,n)$ denote the number of partitions of $n$ with crank $m$, except for [^1] The correct combinatorial values for the anomalous case $n=1$ are $1$ if $m=0$ and 00 for $m\neq 0$. $n=1$ where $M(-1,1)=-M(0,1)=M(1,1):=1$ as given by the following generating function [4]
$C(\zeta;q):=\sum_{\begin{subarray}{c}m\in\mathbb{Z}\\ n\geq 0\end{subarray}}M(m,n)\zeta^{m}q^{n}=\frac{(q)_{\infty}}{(\zeta q)_{\infty}\left(\zeta^{-1}q\right)_{\infty}}.$ (1.2)
Modularity properties of the rank and of the crank generating function differ significantly: The crank generating function is basically a meromorphic Jacobi form, whereas the rank generating function is a “mock Jacobi form” [8]: it only transforms like a Jacobi form after adding a non-holomorphic term (see Section 2.3 for the precise transformations of the rank generating function). We next consider the crank moments [3]
$C_{k}(q):=\sum_{n\geq 0}\sum_{m\in\mathbb{Z}}m^{k}M(m,n)q^{n}.$
These moments can be expressed in terms of quasimodular forms. In [1], the authors expressed these as a so-called partition Eisenstein trace. Here, for a sequence of functions $h=\{h_{k}\}_{k\in\mathbb{N}}$, define, for $n\in\mathbb{N}$, the $n$-th partition trace with respect to $h$ and a function $\phi$ on partitions as
$\mathrm{Tr}_{n}(\phi,h;\tau):=\sum_{\lambda\vdash n}\phi(\lambda)h_{\lambda}(\tau),$
where the sum ranges over all partitions of $n$ and for $\lambda=(1^{m_{1}},2^{m_{2}},\ldots,n^{m_{n}})\vdash n$, we set
$h_{\lambda}(\tau):=\prod_{j=1}^{n}h_{j}^{m_{j}}(\tau).$
Partition Eisenstein traces are the partition traces with respect to the sequence of Eisenstein series $G=\{G_{k}\}_{k\in\mathbb{N}}$. Here, $G_{k}$ is the Eisenstein series [^2] In [1] the authors wrote $G_{k}$ for what here is $2G_{k}$ and $\mathrm{Tr}_{k}(\phi;\tau)$ for what is $\mathrm{Tr}_{2k}(\phi,G;\tau)$ here. of weight $k\in 2\mathbb{N}$, given by
$G_{k}(\tau):=-\frac{B_{k}}{2k}+\sum_{m,n\geq 1}m^{k-1}q^{mn}\qquad(q:=e^{2\pi i\tau},\,\tau\in\mathbb{H}:=\{w\in\mathbb{C}:\operatorname{Im}(w)>0\}),$
with $B_{k}$ the $k$-th Bernoulli number. By convention, $G_{k}:=0$ for $k$ odd. Moreover, define
$\phi(\lambda):=\prod_{j=1}^{k}\frac{2^{m_{j}}}{m_{j}!j!^{m_{j}}}.$ (1.3)
The following result was obtained in [1, Theorem 1.2].
Theorem 1.1.
We have
$\sum_{k\geq 0}C_{k}(q)\frac{z^{k}}{k!}$ $=\frac{2\sinh\!\left(\frac{z}{2}\right)}{z(q)_{\infty}}\sum_{k\geq 0}\mathrm{Tr}_{k}(\phi,G;\tau)z^{k}.$
The Eisenstein series $G_{k}$ play a key role in the theory of modular forms and satisfy many interesting properties. They have a beautiful connection with Bernoulli numbers in that, for $k\geq 2$, we have
$\lim_{\tau\to i\infty}G_{k}(\tau)=-\frac{B_{k}}{2k}.$
For $k>2$, $G_{k}$ is a modular form of weight $k$ on $\mathrm{SL}_{2}(\mathbb{Z})$ and $G_{2}$ is quasimodular (see Subsection 2.1 for the definition). More precisely, we have for $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\mathrm{SL}_{2}(\mathbb{Z})$,
$G_{k}\left(\frac{a\tau+b}{c\tau+d}\right)=\begin{cases}(c\tau+d)^{k}G_{k}(\tau)&\text{if $k\neq 2$},\\ (c\tau+d)^{2}G_{2}(\tau)+\frac{ic}{4\pi}(c\tau+d)&\text{if $k=2$}.\end{cases}$ (1.4)
Another key property of the Eisenstein series is that the corresponding algebra $\mathbb{Q}[G_{2},G_{4},G_{6},\ldots]$ of quasimodular forms is closed under the action of $D:=q\frac{\partial}{\partial q}$.
It is natural to wonder whether a result like Theorem 1.1 involving the rank moments
$R_{k}(q):=\sum_{n\geq 0}\sum_{m\in\mathbb{Z}}m^{k}N(m,n)q^{n}.$ (1.5)
exists. In this paper, we show that this is indeed the case. As mentioned above, the rank moments are related to mock modular forms. Hence, they do not admit an expansion as a trace of modular Eisenstein series like crank moments. As our first result, we write the rank moments as traces of mock Eisenstein series $f_{k}$ defined in Subsection 3.1. The aim of this paper is to understand these functions $f_{k}$. We call a real-analytic function $f^{*}(\tau,\overline{\tau})$ a quasi-completion of $f(\tau)$, if $f^{*}(\tau,\overline{\tau})$ transforms like a quasimodular form (see (2.2)) and if [^3] Here and throughout, we consider $\tau$ and $\overline{\tau}$ as independent variables. $\lim_{\overline{\tau}\to-i\infty}$ $f^{*}(\tau,\overline{\tau})=f(\tau)$. If a quasi-completion $f^{*}(\tau,\overline{\tau})$ transforms as a modular form, then we call $f^{*}(\tau,\overline{\tau})$ a completion. If it is clear from the context, then we also just write $f^{*}(\tau)$ instead of $f^{*}(\tau,\overline{\tau})$. Our first result is the following theorem.
Theorem 1.2.
There exists a family of functions $f=\{f_{k}\}_{k\in\mathbb{N}}$ such that
$\sum_{k\geq 0}R_{k}(q)\frac{z^{k}}{k!}$ $=\frac{2\sinh\!\left(\frac{z}{2}\right)}{z(q)_{\infty}}\sum_{k\geq 0}\mathrm{Tr}_{k}(\phi,f;\tau)z^{k},$ (1.6)
where $\phi$ is defined in (1.3) and $f_{k}$ has the following properties:
  1. (1)
    For $k\geq 2$, we have
    $\lim_{\tau\to i\infty}f_{k}(\tau)=-\frac{B_{k}}{2k}.$
  2. (2)
    The function $f_{k}$ has a quasi-completion [^4] We consider quasi-completions, because they are more analogous to the functions in Theorem 1.1. In the sequel, we also study the completions $\widehat{f_{k}}$ of $f_{k}$. $f^{*}_{k}$ which satisfies, for $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\mathrm{SL}_{2}(\mathbb{Z})$,
    $f^{*}_{k}\left(\frac{a\tau+b}{c\tau+d}\right)=\begin{cases}(c\tau+d)^{k}f^{*}_{k}(\tau)&\text{if $k\neq 2$},\\ (c\tau+d)^{2}f^{*}_{2}(\tau)+\frac{3ic}{4\pi}(c\tau+d)&\text{if $k=2$}.\end{cases}$
  3. (3)
    The algebra $\mathcal{F}:=\mathbb{Q}[f_{2},f_{4},\ldots,G_{2},G_{4},\ldots]$ is closed under the action of $D$.
Remark 1.3.
(1) As for the Eisenstein series, we have $f_{k}=0$ if $k$ is odd; see Remark 3.1.
(2) The transformation of $f_{k}^{*}$ agrees with that of $G_{k}$ up to a factor of $3$ in front of $c\tau+d$. This is explained by the (mock) Jacobi forms underlying the rank and crank statistics: namely, for the rank the corresponding index is $-\frac{3}{2}$ and for the crank it is $-\frac{1}{2}$.
Even though the $f_{k}$ are uniquely determined in Theorem 1.2, properties (1), (2), and (3) itself do not determine them uniquely. For example, after adding a cusp form of weight $k$ to each $f_{k}$, the resulting functions still satisfy these properties. To describe the $f_{k}$ uniquely without involving (1.6), we give two recursive definitions for them involving divisor-like sums. For this, we define $g_{\ell}\in\mathcal{F}$ by $g_{0}:=1$, $g_{\ell}:=0$ for $\ell\in\mathbb{N}$ odd, and, for $\ell\in\mathbb{N}$ even,
$g_{\ell}(\tau):=\left(1-2^{\ell-1}\right)\frac{B_{\ell}}{2\ell}+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\end{subarray}}(2n-3m)^{\ell-1}q^{mn}-\sum_{\begin{subarray}{c}n-1\geq 6m\geq 6\end{subarray}}(n-6m)^{\ell-1}q^{mn}.$ (1.7)
For $r\in\mathbb{N}$, the $r$-th Fourier coefficient of $g_{\ell}$ is a polynomial in some of the divisors of $r$: namely those positive divisors $m,n$ satisfying the inequality $2n-1\geq 3m$ or $n-1\geq 6m$. Hence, our functions are in spirit of the “mock Eisenstein series” studied by Zagier [18, p.15] and Mertens–Ono–Rolen [13, equation (1.4)].
Theorem 1.4.
Let $n\in\mathbb{N}$.
  1. (1)
    We have
    $f_{n}(\tau)$ $=\frac{n}{2^{n-1}}g_{n}(\tau)-\sum_{\ell=2}^{n-2}\frac{\ell}{2^{\ell-2}}\binom{n-1}{\ell}f_{n-\ell}(\tau)g_{\ell}(\tau).$
  2. (2)
    We have
    $\ f_{n}(\tau)$ $=\sum_{\ell=2}^{n}\frac{(n-1)!\ell}{(\ell-1)!2^{\ell-1}}g_{\ell}(\tau)\mathrm{Tr}_{n-\ell}(\psi,f;\tau),$
    with $f:=\{f_{k}(\tau)\}_{k\in\mathbb{N}}$ and $\psi(\lambda):=(-1)^{\sum_{j=1}^{k}m_{j}}\phi(\lambda)$.
The recursive formulas in Theorem 1.4 suggest that the Fourier coefficients of $f_{k}$ may have large denominators. However, we show in the following theorem that, with the exception of the constant term, all Fourier coefficients of $f_{k}$ are integers, similar as for $G_{k}$.
Theorem 1.5.
For $k\geq 2$, the Fourier coefficients of $f_{k}+\frac{B_{k}}{2k}$ are integers.
Finally, in part (1) of the next theorem, we give an explicit formula for $D(f_{k})$ and explain how the raising and lowering operators act on the completions $\widehat{f}_{k}$ of the [^5] See Subsection 2.1 for the definition; in particular, $\widehat{f}_{k}$ and $f_{k}^{*}$ agree if $k\neq 2$. $f_{k}$. More precisely, let ^$\widehat{\mathcal{F}}:=\mathbb{C}[\widehat{f}_{2},\widehat{f}_{4},\widehat{f}_{6},\ldots,\widehat{G}_{2},G_{4},G_{6}]$ and define the raising and the lowering operator by
$\mathcal{R}_{k}:=2i\frac{\partial}{\partial\tau}+\frac{k}{v},\qquad L:=-2iv^{2}\frac{\partial}{\partial\overline{\tau}}\qquad\qquad(k\in\mathbb{Z},\tau=u+iv).$
In parts (2) and (3), we obtain a recursive expression for the action of the raising and lowering operators on $\widehat{f}_{k}$. In the physics literature, the latter of these falls into the realm of holomorphic anomaly equations. We let $\delta_{\mathcal{S}}:=1$ if a statement $\mathcal{S}$ holds and $\delta_{\mathcal{S}}:=0$ otherwise.
Theorem 1.6.
We have the following.
  1. (1)
    For $k\geq 2$, we have
    $D(f_{k}(\tau))=\frac{k!}{6}\mathrm{Tr}_{k+2}(\phi,3G-f;\tau)-\frac{k-1}{6(k+1)}f_{k+2}(\tau)-\frac{1}{3}\sum_{a=1}^{k-1}\binom{k}{a}f_{a+1}(\tau)f_{k-a+1}(\tau).$
  2. (2)
    The algebra $\widehat{\mathcal{F}}$ is closed under the raising operator. In particular, for $k\geq 2$ we have
    $\hskip 14.22636pt-\frac{1}{4\pi}\mathcal{R}_{k}\!\left(\widehat{f_{k}}(\tau)\right)=\frac{k!}{6}\mathrm{Tr}_{k+2}\left(\phi,3\widehat{G}-\widehat{f};\tau\right)-\frac{k-1}{6(k+1)}\widehat{f}_{k+2}(\tau)\\ -\frac{1}{3}\sum_{a=1}^{k-1}\binom{k}{a}\widehat{f}_{a+1}(\tau)\widehat{f}_{k-a+1}(\tau),$
    where $\widehat{G}=\{\widehat{G}_{k}\}_{k\geq 1}$ is defined by $\widehat{G}_{k}=G_{k}$ for $k\neq 2$ and $\widehat{G}_{2}$ is the completion of $G_{2}$, given in (2.1).
  3. (3)
    We have $L(\widehat{\mathcal{F}})\subseteq\widehat{\mathcal{F}}\oplus\sqrt{v}|\eta|^{2}\widehat{\mathcal{F}}$. In particular, for $k\geq 2$ we have
    $L\!\left(\widehat{f}_{k}(\tau)\right)$ $=-\frac{3}{8\pi}\delta_{k=2}+\frac{\sqrt{3}k!}{4\sqrt{2}\pi}\sqrt{v}|\eta(\tau)|^{2}\mathrm{Tr}_{k-2}\left(\psi,\widehat{f};\tau\right),$
    where $\psi$ is defined in Theorem 1.4 (2).
  4. (4)
    If $\widehat{f}\in\widehat{\mathcal{F}}$ is of weight $k$, then $(D+\frac{2k}{3}f_{2}^{*})(\widehat{f})$ transforms modular of weight $k+2$.
Remark 1.7.
We expect that the algebra $\mathcal{F}$ is freely generated by the $f_{k}$ for $k$ even and $G_{2},G_{4}$, and $G_{6}$ (see also Question (2) in Section 6). In particular, this would imply that the subalgebra $\mathbb{Q}[f_{2},f_{4},\ldots]$ of $\mathcal{F}$ is not closed under the derivation $D$, since Theorem 1.6 (1) involves the modular Eisenstein series.
The paper is organized as follows. In Section 2, we provide certain preliminaries on quasimodular forms and crank moments, the completion of the rank generating function, the rank-crank PDE, Pólya cycle index polynomials, and finally, divisibility properties of multinomials. In Section 3 and Section 4, we prove our theorems. The last two sections are devoted to examples and open questions.
Acknowledgements
We thank Caner Nazaroglu and Kilian Rausch for helpful comments on a previous version of the manuscript. The first and the second authors were funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001179) and the first and the third authors by the SFB/TRR 191 “Symplectic Structure in Geometry, Algebra and Dynamics”, funded by the DFG (Projektnummer 281071066 TRR 191).
2. Preliminaries
2.1. Modular forms and quasimodular forms
The Eisenstein series $G_{k}$ are modular forms of weight $k$ for $\mathrm{SL}_{2}(\mathbb{Z})$ for $k\geq 4$ even. If $k=2$, then we need to add a non-holomorphic part to make $G_{2}$ modular. To be more precise,
$\widehat{G}_{2}(\tau):=G_{2}(\tau)+\frac{1}{8\pi v}$ (2.1)
transforms like a modular form of weight $2$. The holomorphic part $G_{2}$ can be recovered by
$\lim_{\overline{\tau}\to-i\infty}\widehat{G}_{2}(\tau)=G_{2}(\tau).$
In general $f:\mathbb{H}\to\mathbb{C}$ is an almost holomorphic modular form of weight $k\in\mathbb{Z}$ and depth $s\in\mathbb{N}_{0}$, if the following conditions hold:
  1. (1)
    We have, for $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$,
    $f\left(\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{k}f(\tau).$
  2. (2)
    We have [^6] The $f_{j}$ here should not be confused with the “mock Eisenstein series” $f_{j}$.
    $f(\tau)=\sum_{j=0}^{s}\frac{f_{j}(\tau)}{v^{j}},$
    for some holomorphic functions $f_{j}:\mathbb{H}\to\mathbb{C}$ with $f_{s}\neq 0$.
  3. (3)
    The function $f$ grows at most polynomially in $\frac{1}{v}$ as $v\to 0$.
By convention, the zero function is an almost holomorphic modular form of depth $\infty$- ∞.
The holomorphic part $f_{0}$ is called a quasimodular form of weight $k$ and depth $s$. Note that
$f_{0}(\tau)=\lim_{\overline{\tau}\to-i\infty}f(\tau).$
More concretely, for a quasimodular form $g$ of weight $k$ and depth $s$ there exist holomorphic functions $g_{j}:\mathbb{H}\to\mathbb{C}$ for $j\in\{0,\ldots,s\}$ with $g_{0}=g$ such that
$(c\tau+d)^{-k}g\!\left(\frac{a\tau+b}{c\tau+d}\right)\>=\>\sum_{j=0}^{s}g_{j}(\tau)\Bigl{(}\frac{c}{c\tau+d}\Bigr{)}^{j}$
for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$. Similarly, we say an analytic function $g^{*}(\tau,\overline{\tau})$ transforms like a quasimodular form if there exist real-analytic functions $g_{j}^{*}(\tau,\overline{\tau})$ so that
$(c\tau+d)^{-k}g^{*}\!\left(\frac{a\tau+b}{c\tau+d},\frac{a\overline{\tau}+b}{c\overline{\tau}+d}\right)\>=\>\sum_{j=0}^{s}g_{j}(\tau,\overline{\tau})\Bigl{(}\frac{c}{c\tau+d}\Bigr{)}^{j}$ (2.2)
for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$. The space of all quasimodular forms is a free algebra with generators $G_{2},G_{4}$, and $G_{6}$. It follows from Ramanujan’s differential equations, [14, equations (1), (2)]
$D(G_{2})$ $=-2G_{2}^{2}+\frac{5}{6}G_{4},\quad D(G_{4})=-8G_{2}G_{4}+\frac{7}{10}G_{6},\quad D(G_{6})=-12G_{2}G_{6}+\frac{400}{7}G_{4}^{2},$ (2.3)
that this algebra is closed under differentiation. We also require the Serre derivative (see, e.g., [17, p. 48])
$\vartheta_{k}:=D+2kG_{2}$ (2.4)
which acts on modular forms of weight $k$, preserving the algebra of modular forms.
The Dedekind eta function
$\eta(\tau):=q^{\frac{1}{24}}\prod_{n\geq 1}\left(1-q^{n}\right)$
is a modular form of weight $\frac{1}{2}$. It is not hard to deduce the following lemma.
Lemma 2.1.
For any holomorphic function $f$ on $\mathbb{H}$, we have
$\eta D\!\left(\frac{f}{\eta}\right)=G_{2}f+D(f).$
2.2. Crank moments
Set $\zeta:=e^{2\pi iz}$. By (1.2) and [16, equation (7)], the crank generating function can be expressed in terms of the Eisenstein series. Recall that $G_{k}=0$ if $k$ is odd.
Lemma 2.2.
We have
$C(\zeta;q)=\frac{\sin(\pi z)}{\pi z(q)_{\infty}}\exp\!\left(2\sum_{k\geq 2}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right).$
Taking $\tau\to i\infty$, we obtain the following lemma, which was also observed in [2, Lemma 3.1].
Lemma 2.3.
We have
$\frac{\zeta^{\frac{1}{2}}}{\zeta-1}=\frac{1}{2\pi iz}\exp\!\left(-\sum_{k\geq 2}\frac{B_{k}}{k}\frac{(2\pi iz)^{k}}{k!}\right).$
Recall that the Bernoulli numbers $B_{n}$ are defined as the constant terms of the Bernoulli polynomials $B_{n}(x)$ of degree $n$, which satisfy
$\sum_{n\geq 0}B_{n}(X){\frac{t^{n}}{n!}}$ $={\frac{te^{Xt}}{e^{t}-1}},$ (2.5)
$B_{n}(X+Y)$ $=\sum_{k=0}^{n}{n\choose k}B_{n-k}(X)Y^{k},$ (2.6)
$B_{n}^{\prime}(X)$ $=nB_{n-1}(X).$ (2.7)
2.3. Mock modularity of the rank generating function
The rank generating function can be written as a Lerch sum
$R(\zeta;q)=\frac{1-\zeta}{(q)_{\infty}}\sum_{n\in\mathbb{Z}}\frac{(-1)^{n}q^{\frac{n(3n+1)}{2}}}{1-\zeta q^{n}}.$ (2.8)
Note that this closely resembles the following representation of the crank generating function
$C(\zeta;q)=\frac{1-\zeta}{(q)_{\infty}}\sum_{n\in\mathbb{Z}}\frac{(-1)^{n}q^{\frac{n(n+1)}{2}}}{1-\zeta q^{n}}.$
In Subsection 4.1 of [6], the first author defined (using different notation)
$R^{\#}(z;\tau):=\left(\frac{R(\zeta;q)}{\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}}q^{-\frac{1}{24}}+\frac{1}{2}q^{-\frac{1}{6}}\sum_{\pm}\pm\zeta^{\mp 1}S(3z\pm\tau;3\tau)\right)e^{12\pi^{2}G_{2}(\tau)z^{2}},$ (2.9)
where ($\tau=u+iv,z=x+iy,u,v,x,y\in\mathbb{R}$)
$S(z;\tau):=\sum_{n\in\mathbb{Z}+\frac{1}{2}}\left(\operatorname{sgn}(n)-E\left(\left(n+\frac{y}{v}\right)\sqrt{2v}\right)\right)(-1)^{n-\frac{1}{2}}q^{-\frac{n^{2}}{2}}e^{-2\pi inz},$
[…]
Lemma 2.4.
For $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$, we have
$\eta\!\left(\frac{a\tau+b}{c\tau+d}\right)R^{\#}\!\left(\frac{z}{c\tau+d};\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)\eta(\tau)R^{\#}(z;\tau).$
2.4. The rank-crank PDE
The rank-crank PDE of Atkin and Garvan [3, Theorem 1.1] relates the rank and crank generating functions by a differential equation. To state it, we define the heat operator [^7] We point out that we re-normalize the standard heat operator by multiplying by a factor of $\frac{1}{4\pi^{2}}$ (see [11, p. 33] for more details). (of index $-\frac{3}{2}$) as
$H:=6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{\!2}.$
This heat operator maps Jacobi forms of weight $\frac{1}{2}$ to weight $\frac{5}{2}$ and does not change the index. We state a modified version of the rank-crank PDE which is more convenient for us.
Lemma 2.5.
We have
$2\left(\frac{\zeta^{\frac{1}{2}}(q)_{\infty}C(\zeta;q)}{1-\zeta}\right)^{\!3}=(H+6G_{2}(\tau))\left(\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R(\zeta;q)}{1-\zeta}\right).$
Proof.
In [7, Theorem 14.28], the rank-crank PDE of Atkin–Garvan was formulated as
$2\eta^{2}(\tau)\left(\frac{\zeta^{\frac{1}{2}}q^{-\frac{1}{24}}C(\zeta;q)}{1-\zeta}\right)^{\!3}=H\left(\frac{\zeta^{\frac{1}{2}}q^{-\frac{1}{24}}R(\zeta;q)}{1-\zeta}\right).$
Multiplying both sides by $\eta(\tau)=q^{\frac{1}{24}}(q)_{\infty}$ yields
$2\left(\frac{\zeta^{\frac{1}{2}}(q)_{\infty}C(\zeta;q)}{1-\zeta}\right)^{\!3}=\eta(\tau)H\left(\frac{1}{\eta(\tau)}\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R(\zeta;q)}{1-\zeta}\right).$
Using the definition of $H$ and Lemma 2.1 with
$f(z;\tau)=\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R(\zeta;q)}{1-\zeta}$
we obtain the lemma. ∎
We also require the following lemma, inspired by the observation in [9] that the non-holomorphic part of a certain non-holomorphic Jacobi form, closely related to $R^{\#}$, is annihilated by the heat operator.
Lemma 2.6.
We have
$H\!\left(R^{\#}(z;\tau)e^{-12\pi^{2}G_{2}(\tau)z^{2}}\right)=-H\!\left(\frac{\zeta^{\frac{1}{2}}R(\zeta;q)q^{-\frac{1}{24}}}{1-\zeta}\right).$
Proof.
Using (2.9), it is enough to show
$H\!\left(q^{-\frac{1}{6}}\zeta^{-1}S(3z+\tau;3\tau)\right)=0.$
From (4.1) of [6], we have
$q^{-\frac{1}{6}}\zeta^{-1}S(3z+\tau;3\tau)=\!\sum_{n\in\mathbb{Z}-\frac{1}{6}}\!\left(\operatorname{sgn}\!\left(n-\frac{1}{3}\right)-E\!\left(\left(n+\frac{y}{v}\right)\sqrt{6v}\right)\!\right)(-1)^{n-\frac{5}{6}}q^{-\frac{3n^{2}}{2}}\zeta^{-3n}.$ (2.10)
Thus the claim follows once we show that
$H\!\left(\!\left(\operatorname{sgn}\left(n-\frac{1}{3}\right)-E\left(\left(n+\frac{y}{v}\right)\sqrt{6v}\right)\right)q^{-\frac{3n^{2}}{2}}\zeta^{-3n}\right)=0.$
Noting that $H(q^{-\frac{3n^{2}}{2}}\zeta^{-3n})=0$ and
$H(fg)=H(f)g+fH(g)+2\left(\zeta\frac{\partial}{\partial\zeta}f\right)\left(\zeta\frac{\partial}{\partial\zeta}g\right),$
it is enough to prove that
$0=H\!\left(E\left(\left(n\!+\!\frac{y}{v}\right)\sqrt{6v}\right)\right)-\frac{3n}{\pi i}\frac{\partial}{\partial z}E\left(\left(n\!+\!\frac{y}{v}\right)\sqrt{6v}\right)=\frac{3}{8\pi^{2}v}[2\pi wE^{\prime}(w)+E^{\prime\prime}(w)]_{w=\left(n+\frac{y}{v}\right)\sqrt{6v}}\raisebox{-5.0pt}{.}$
Thus we want to show that
$2\pi xE^{\prime}(x)+E^{\prime\prime}(x)=0,$
which holds since $E^{\prime}(x)=2e^{-\pi x^{2}}$ and $E^{\prime\prime}(x)=-4\pi xe^{-\pi x^{2}}$. ∎
2.5. Pólya cycle index polynomials
We require a result about Pólya cycle index polynomials in the case of symmetric group $S_{n}$, the set of permutations of the symbols $x_{1},x_{2},\ldots,x_{n}$; see [15] and [1, Lemma 2.1] for more details.
Lemma 2.7 (Example 5.2.10 of [15]).
We have
$\sum_{{n\geq 0}}\sum_{\lambda\vdash n}\prod_{k=1}^{n}\frac{x_{k}^{m_{k}}}{m_{k}!}w^{n}=\exp\left(\sum_{{k\geq 1}}x_{k}w^{k}\right)$
as a formal power series in $w$, where $\lambda=(1^{m_{1}},2^{m_{2}},\ldots,n^{m_{n}})\vdash n$.
2.6. Divisibility of multinomials
Below, we require the following lemma regarding the divisibility of multinomial coefficients, which follows directly from Bézout’s lemma.
Lemma 2.8.
For positive integers $a_{1},\ldots,a_{\ell}$ with sum $\sum_{j=1}^{\ell}a_{j}=n$, we have
$\frac{n}{\gcd(a_{1},\ldots,a_{\ell})}\,\Big{|}\,\binom{n}{a_{1},a_{2},\ldots,a_{\ell}}.$
3. Mock Eisenstein series and the proof of Theorems 1.2 and 1.6
3.1. Modularity and completion
Motivated by Lemma 2.2, we define the mock Eisenstein series $f_{k}$ in terms of the rank generating function as follows:
$R(\zeta;q)=:\frac{\sin\!\left(\pi z\right)}{\pi z(q)_{\infty}}\exp\!\left(2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right).$ (3.1)
Remark 3.1.
Since $R$ is invariant under $\zeta\mapsto\zeta^{-1}$, i.e., $z\mapsto-z$, we have $f_{k}=0$ if $k$ is odd.
In the remainder of this section, we show that the $f_{k}$ satisfies all of the properties in Theorem 1.2. To define their quasi-completions, we let
$R^{\circ}(z,\overline{z};\tau,\overline{\tau}):=$ $2i\sin(\pi z)q^{\frac{1}{24}}R^{\#}(z;\tau)e^{-12\pi^{2}G_{2}(\tau)z^{2}}$
=\ = $R(\zeta;q)+\frac{1}{2}q^{-\frac{1}{8}}\left(\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}\right)\sum_{\pm}\pm\zeta^{\mp 1}S(3z\pm\tau;3\tau).$ (3.2)
Lemma 3.2.
We have
$\lim_{\overline{\tau}\to-i\infty}R^{\circ}(z,\overline{z};\tau,\overline{\tau})=R(\zeta;q).$
Proof.
By (3.2), we have
$\lim_{\overline{\tau}\to-i\infty}R^{\circ}(z,\overline{z};\tau,\overline{\tau})=R(\zeta;q)+\frac{1}{2}q^{-\frac{1}{8}}\left(\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}\right)\sum_{\pm}\pm\zeta^{\mp 1}\lim_{\overline{\tau}\to-i\infty}S(3z\pm\tau;3\tau).$
From equation (4.1) of [6], we directly obtain
$\lim_{\overline{\tau}\to-i\infty}q^{-\frac{1}{6}}\zeta^{-1}S(3z+\tau;3\tau)\\ =\sum_{n\in\mathbb{Z}-\frac{1}{6}}\left(\operatorname{sgn}\left(n-\frac{1}{3}\right)-\lim_{\overline{\tau}\to-i\infty}E\left(\left(n+\frac{y}{v}\right)\sqrt{6v}\right)\right)(-1)^{n-\frac{5}{6}}q^{-\frac{3n^{2}}{2}}\zeta^{-3n}.$
Now
$\lim_{\overline{\tau}\to-i\infty}E\left(\left(n+\frac{y}{v}\right)\sqrt{6v}\right)=\lim_{v\to\infty}E\left(n\sqrt{6v}+\frac{\sqrt{6}y}{\sqrt{v}}\right)=\operatorname{sgn}(n).$
As $\operatorname{sgn}(n-\frac{1}{3})=\operatorname{sgn}(n)$ for $n\in\mathbb{Z}-\frac{1}{6}$ and $S(3z-\tau;3\tau)=S(-3z+\tau;3\tau)$, the claimed statement follows. ∎
We next introduce quasi-completions $f_{k}^{*}$ of $f_{k}$. First, we define
$\frac{(q)_{\infty}R^{\circ}(z,\overline{z};\tau,\overline{\tau})}{2i\sin(\pi z)}=:\frac{1}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}(\tau,\overline{\tau})\frac{(2\pi iz)^{k}}{k!}\frac{(2\pi i\overline{z})^{\ell}}{\ell!}\right).$ (3.3)
We let [^8] Note that here we suppress the dependence on $\overline{\tau}$. $R^{*}(z;\tau)$ be the constant term of $R^{\circ}(z,\overline{z};\tau,\overline{\tau})$ in the Taylor expansion in $\overline{z}$ and set
$f^{*}_{k}(\tau):=f^{*}_{k,0}(\tau,\overline{\tau}).$
In terms of these, we define [^9] Again $\mathbb{F}(z;\tau)$ depends on $\overline{\tau}$.
$\mathbb{F}(z;\tau):=\frac{(q)_{\infty}R^{*}(z;\tau)}{2i\sin(\pi z)}=\frac{1}{2\pi iz}\exp\!\left(2\sum_{k\geq 0}f^{*}_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right).$ (3.4)
Note that $f^{*}_{k}=0$ for $k$ odd. We are now ready to prove Theorem 1.2 (2).
Lemma 3.3.
The function $f_{k}^{*}$ is a quasi-completion of $f_{k}$. In particular, for $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\mathrm{SL}_{2}(\mathbb{Z})$,
$f^{*}_{k}\left(\frac{a\tau+b}{c\tau+d}\right)=\begin{cases}(c\tau+d)^{k}f^{*}_{k}(\tau)&\text{if $k\neq 2$},\\ (c\tau+d)^{2}f^{*}_{2}(\tau)+\frac{3ic}{4\pi}(c\tau+d)&\text{if $k=2$}.\end{cases}$
Proof.
We start by showing the transformation law of $f^{*}_{k}$. Combining (3.3) and (3.2) with (1.4) and Lemma 2.4, we obtain
$\frac{c\tau+d}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}\left(\frac{a\tau+b}{c\tau+d},\frac{a\overline{\tau}+b}{c\overline{\tau}+d}\right)\frac{\left(\frac{2\pi iz}{c\tau+d}\right)^{k}}{k!}\frac{\left(\frac{2\pi i\overline{z}}{c\overline{\tau}+d}\right)^{\ell}}{\ell!}\right)$
$\hskip 56.9055pt=\eta\left(\frac{a\tau+b}{c\tau+d}\right)R^{\#}\left(\frac{z}{c\tau+d};\frac{a\tau+b}{c\tau+d}\right)e^{-12\pi^{2}G_{2}\left(\frac{a\tau+b}{c\tau+d}\right)\left(\frac{z}{c\tau+d}\right)^{2}}$
$\hskip 56.9055pt=(c\tau+d)\eta(\tau)R^{\#}(z;\tau)e^{-12\pi^{2}\left(G_{2}(\tau)+\frac{ic}{4\pi(c\tau+d)}\right)z^{2}}$
$\hskip 56.9055pt=\frac{c\tau+d}{2\pi iz}\exp\!\left(2\left(\frac{3ic}{4\pi(c\tau+d)}\frac{(2\pi iz)^{2}}{2!}+\sum_{k,\ell\geq 0}f^{*}_{k,\ell}(\tau,\overline{\tau})\frac{(2\pi iz)^{k}}{k!}\frac{(2\pi i\overline{z})^{\ell}}{\ell!}\right)\right).$
Hence, we get
$f^{*}_{k,\ell}\left(\frac{a\tau+b}{c\tau+d},\frac{a\overline{\tau}+b}{c\overline{\tau}+d}\right)=\begin{cases}(c\tau+d)^{k}(c\overline{\tau}+d)^{\ell}f^{*}_{k,\ell}(\tau,\overline{\tau})&\text{if }(k,\ell)\neq(2,0),\\ (c\tau+d)^{2}f^{*}_{2,0}(\tau,\overline{\tau})+\frac{3ic}{4\pi}(c\tau+d)&\text{if }(k,\ell)=(2,0).\end{cases}$
So the transformation formula for $f^{*}_{k}$ follows.
We next show that
$\lim_{\overline{\tau}\to-i\infty}f^{*}_{k}(\tau)=f_{k}(\tau).$ (3.5)
By (3.3), Lemma 3.2, and (3.1) we have
$\frac{1}{2\pi iz}\lim_{\overline{\tau}\to-i\infty}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}(\tau,\overline{\tau})\frac{(2\pi iz)^{k}}{k!}\frac{(2\pi i\overline{z})^{\ell}}{\ell!}\right)=\frac{(q)_{\infty}}{2i\sin(\pi z)}\lim_{\overline{\tau}\to-i\infty}R^{\circ}(z,\overline{z};\tau,\overline{\tau})\hskip 14.22636pt$
$=\frac{(q)_{\infty}}{2i\sin(\pi z)}R(\zeta;q)=\frac{1}{2\pi iz}\exp\!\left(2\sum_{k\geq 0}f_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right).$
Comparing coefficients gives (3.5). ∎
We also define
$\widehat{\mathbb{F}}(z;\tau):=\frac{(q)_{\infty}R^{*}(z;\tau)e^{-\frac{3\pi z^{2}}{2v}}}{2i\sin(\pi z)}=:\frac{1}{2\pi iz}\exp\!\left(2\sum_{k\geq 1}\widehat{f}_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right).$ (3.6)
Then, by the same argument used to prove Lemma 3.3, we obtain the following result.
Lemma 3.4.
For $k\in\mathbb{N}$, we have
$\widehat{f}_{k}(\tau)=\begin{cases}f_{k}^{*}(\tau)&\text{if $k\neq 2$},\\ f_{2}^{*}(\tau)+\frac{3}{8\pi v}&\text{if $k=2$}.\end{cases}$
In particular
$\widehat{f}_{k}\left(\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{k}\widehat{f}_{k}(\tau).$
3.2. Limiting behavior of $f_{k}$
Next, we determine the behavior of $f_{k}(\tau)$ as $\tau\to i\infty$.
Lemma 3.5.
For $k\geq 2$, we have
$\lim_{\tau\to i\infty}f_{k}(\tau)=\lim_{\tau\to i\infty}f^{*}_{k}(\tau)=\lim_{\tau\to i\infty}\widehat{f}_{k}(\tau)=-\frac{B_{k}}{2k}.$
Proof.
Using (2.8), we first compute
$\frac{(q)_{\infty}R(\zeta;q)}{2i\sin(\pi z)}=-\frac{\zeta^{\frac{1}{2}}(q)_{\infty}}{1-\zeta}R(\zeta;q)=-\zeta^{\frac{1}{2}}\sum_{n\in\mathbb{Z}}\frac{(-1)^{n}q^{\frac{n(3n+1)}{2}}}{1-\zeta q^{n}}\to-\frac{\zeta^{\frac{1}{2}}}{1-\zeta}$
as $\tau\to i\infty$. Thus, by (3.1),
$\frac{\zeta^{\frac{1}{2}}}{\zeta-1}=\frac{1}{2\pi iz}\exp\!\left(2\sum_{k\geq 1}\lim_{\tau\to i\infty}f_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right).$
By Lemma 2.3, we obtain the claim for $f_{k}$.
To prove the claim for $f^{*}_{k}$, by (3.1) and (3.4), we have to show that
$\lim_{\tau\to i\infty}\frac{(q)_{\infty}\left(R^{*}(z;\tau)-R(\zeta;q)\right)}{2i\sin(\pi z)}=0.$ (3.7)
By (3.2), we have
$R^{*}(z;\tau)-R(\zeta;q)=\frac{1}{2}q^{-\frac{1}{8}}\left(\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}\right)\sum_{\pm}\pm\zeta^{\mp 1}S(3z\pm\tau;3\tau).$
Next, note that
$S(3z\pm\tau;3\tau)=\sum_{n\in\mathbb{Z}+\frac{1}{2}}\left(\operatorname{sgn}(n)-E\left(\left(n\pm\frac{1}{3}+\frac{y}{v}\right)\sqrt{6v}\right)\right)(-1)^{n-\frac{1}{2}}q^{-\frac{3n^{2}}{2}}e^{-2\pi in(3z\pm\tau)}.$
Using Lemma 1.7 of [19], as $\tau\to i\infty$ (so $v\to\infty$),
$\operatorname{sgn}(n)-E\left(\left(n\pm\frac{1}{3}+\frac{y}{v}\right)\sqrt{6v}\right)\ll e^{-\pi\left(n\pm\frac{1}{3}+\frac{y}{v}\right)^{2}6v}.$
Thus $S(3z\pm\tau;3\tau)$ is bounded as $\tau\to i\infty$. Multiplying with $(q)_{\infty}$ and taking the limit gives (3.7). The claim for $\widehat{f}_{k}$ follows from the claim for $f_{k}^{*}$ and Lemma 3.4. ∎
3.3. Differential equations
Next, we look at the action of $D$ on the $f_{k}$ as well as the action of the raising and the lowering operator on its completions.
Proof of Theorem 1.6.
[…]
$\frac{2}{(2\pi iz)^{3}}\exp\!\left(6\sum_{k\geq 1}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right)$
$=\left(6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{2}+6G_{2}(\tau)\right)\frac{1}{2\pi iz}\exp\!\left(2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right)$ (3.8)
$=\left(\!\left(12\sum_{k\geq 1}D(f_{k}(\tau))\frac{(2\pi iz)^{k}}{k!}+6G_{2}(\tau)\right)\frac{1}{2\pi iz}+\frac{2}{(2\pi iz)^{3}}+2\sum_{k\geq 1}f_{k}(\tau)\frac{(k-2)(2\pi iz)^{k-3}}{(k-1)!}\right.$
$\hskip 5.69046pt\left.+2\left(-\frac{1}{(2\pi iz)^{2}}+2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k-2}}{(k-1)!}\right)\sum_{\ell\geq 1}f_{\ell}(\tau)\frac{(2\pi iz)^{\ell-1}}{(\ell-1)!}\right)\exp\!\left(2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right).$
Multiplying by $\frac{(2\pi iz)^{3}}{2}\exp(-2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k}}{k!})$ and collecting all terms with $D(f_{k})$ on one side and the rest on the other, we have
$6\sum_{k\geq 1}D(f_{k}(\tau))\frac{(2\pi iz)^{k+2}}{k!}=-3G_{2}(\tau)(2\pi iz)^{2}+\exp\!\left(2\sum_{k\geq 1}(3G_{k}(\tau)-f_{k}(\tau))\frac{(2\pi iz)^{k}}{k!}\right)-1\\ -\sum_{k\geq 1}f_{k}(\tau)(k-3)\frac{(2\pi iz)^{k}}{(k-1)!}-2\left(\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k}}{(k-1)!}\right)^{2}.$ (3.9)
Using Lemma 2.7 with $w=2\pi iz$ and ${x_{k}}=\frac{2}{k!}(3G_{k}-f_{k})$ we obtain
$\exp\!\left(2\sum_{k\geq 1}(3G_{k}(\tau)-f_{k}(\tau))\frac{(2\pi iz)^{k}}{k!}\right)=\sum_{k\geq 0}\mathrm{Tr}_{k}(\phi,3G-f;\tau)(2\pi iz)^{k}.$
Plugging this in (3.9) and extracting the coefficient of ${(2\pi iz)^{k+2}}$ on both sides, we get, for $k\geq 2$,
$\frac{6}{k!}D(f_{k}(\tau))$ $=\mathrm{Tr}_{k+2}(\phi,3G-f;\tau)-\frac{k-1}{(k+1)!}f_{k+2}(\tau)-2\sum_{\begin{subarray}{c}a,b\geq 0\\ a+b=k\end{subarray}}\frac{1}{a!b!}f_{a+1}(\tau)f_{b+1}(\tau).$
Multiplying with $\frac{k!}{6}$, yields the claim.
(2) By (3.2) and Lemma 2.6, we have
$H\!\left(\frac{1}{\eta(\tau)}\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R^{\circ}(z,\overline{z};\tau,\overline{\tau})}{1-\zeta}\right)=H\!\left(\frac{1}{\eta(\tau)}\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R(\zeta;q)}{1-\zeta}\right).$
Using Lemma 2.1 and then Theorem 2.5, we obtain
$(H+6G_{2}(\tau))\left(\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R^{\circ}(z,\overline{z};\tau,\overline{\tau})}{1-\zeta}\right)=2\left(\frac{\zeta^{\frac{1}{2}}(q)_{\infty}C(\zeta;q)}{1-\zeta}\right)^{\!3}.$
Employing (3.3) for the left-hand side, we obtain
$2\left(\frac{(q)_{\infty}C(\zeta;q)}{2i\sin(\pi z)}\right)^{\!3}=\left(H+6G_{2}(\tau)\right)\left(\frac{1}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}(\tau,\overline{\tau})\frac{(2\pi iz)^{k}}{k!}\frac{(2\pi i\overline{z})^{\ell}}{\ell!}\right)\right).$
Hence, using Lemma 2.2 and the definition of $H$, we obtain
$\frac{2}{(2\pi iz)^{3}}\exp\!\left(6\sum_{k\geq 1}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right)\\ =\left(6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{2}+6G_{2}(\tau)\right)\frac{1}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}(\tau,\overline{\tau})\frac{(2\pi iz)^{k}}{k!}\frac{(2\pi i\overline{z})^{\ell}}{\ell!}\right).$
Taking the constant terms from both sides with respect to $\overline{z}$ gives
$\frac{2}{(2\pi iz)^{3}}\exp\!\left(\!6\sum_{k\geq 1}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\!\right)\\ =\left(\!6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{2}+6G_{2}(\tau)\!\right)\!\frac{1}{2\pi iz}\exp\!\left(\!2\sum_{k\geq 0}f^{*}_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\!\right)\!.$
Now the shape of the above equation is exactly the same as in (3.8) except that $f_{k}$ is replaced by $f^{*}_{k}$ on the right-hand side. Hence, the same calculation as in part (1) gives that
$\frac{6}{k!}D\left(f^{*}_{k}(\tau)\right)=\mathrm{Tr}_{k+2}\left(\phi,3G-f^{*};\tau\right)-\frac{k-1}{(k+1)!}f^{*}_{k+2}(\tau)-2\sum_{\begin{subarray}{c}a,b\geq 0\\ a+b=k\end{subarray}}\frac{1}{a!b!}f^{*}_{a+1}(\tau)f^{*}_{b+1}(\tau).$
Noting that $3G-f^{*}=3\widehat{G}-\widehat{f}$, by Lemma 3.4, we have
$\frac{6}{k!}D\left(f^{*}_{k}(\tau)\right)=\mathrm{Tr}_{k+2}\left(\phi,3\widehat{G}-\widehat{f};\tau\right)-\frac{k-1}{(k+1)!}f^{*}_{k+2}(\tau)-2\sum_{\begin{subarray}{c}a,b\geq 0\\ a+b=k\end{subarray}}\frac{1}{a!b!}f^{*}_{a+1}(\tau)f^{*}_{b+1}(\tau).$
From this, it is not hard to conclude the claim.
(3) Since $\widehat{\mathcal{F}}$ is generated by $\widehat{f}_{k}$ for $k\in\mathbb{N}$, $\widehat{G}_{2},G_{4}$, and $G_{6}$, it is enough to show that $L(\widehat{f}_{k}),L(\widehat{G}_{2})$, $L(G_{4}),L(G_{6})\in\widehat{\mathcal{F}}\oplus\sqrt{v}|\eta|^{2}\widehat{\mathcal{F}}$ and prove the formula for $L(\widehat{f}_{k})$. First, we have $L(G_{4})=L(G_{6})=0\in\widehat{\mathcal{F}}\oplus\sqrt{v}|\eta|^{2}\widehat{\mathcal{F}}$ since $G_{4}$ and $G_{6}$ are holomorphic. Next, by (2.1), we have $L(\widehat{G}_{2})=-\frac{1}{8\pi}\in\widehat{\mathcal{F}}\oplus\sqrt{v}|\eta|^{2}\widehat{\mathcal{F}}$. Finally, we look at $L(\widehat{f}_{k})$. We compute, using (3.6),
$L\!\left(\widehat{\mathbb{F}}(z;\tau)\right)=2\widehat{\mathbb{F}}(z;\tau)\sum_{k\geq 0}L\!\left(\widehat{f}_{k}(\tau)\right)\frac{(2\pi iz)^{k}}{k!}.$ (3.10)
Using (3.6) again, the left-hand side can be written as
$L\!\left(\frac{(q)_{\infty}R^{*}(z;\tau)e^{-\frac{3\pi z^{2}}{2v}}}{2i\sin(\pi z)}\right)=\frac{(q)_{\infty}}{2i\sin(\pi z)}\left(L\!\left(R^{*}(z;\tau)\right)+R^{*}(z;\tau)\frac{3\pi z^{2}}{2}\right)e^{-\frac{3\pi z^{2}}{2v}}.$
Using the above for the left-hand side and (3.6) for the right-hand side of (3.10) gives
$L\!\left(R^{*}(z;\tau)\right)=R^{*}(z;\tau)\sum_{k\geq 0}\left(2L\!\left(\widehat{f}_{k}(\tau)\right)+\frac{3}{4\pi}\delta_{k=2}\right)\frac{(2\pi iz)^{k}}{k!}.$ (3.11)
We next compute, using (3.2),
$L\!\left(R^{*}(z;\tau)\right)=\frac{1}{2}q^{-\frac{1}{8}}\left(\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}\right)\sum_{\pm}\pm\zeta^{\mp 1}L\!\left(S(3z\pm\tau;3\tau)\right).$ (3.12)
We first consider the plus sign. Applying the lowering operator to (2.10) and using that by [6, p. 11], we have
$L\!\left(E\left(\left(n+\frac{y}{v}\right)\sqrt{6v}\right)\right)=\sqrt{6}v^{\frac{3}{2}}\left(n-\frac{y}{v}\right)e^{-6\pi v\left(n+\frac{y}{v}\right)^{2}}$
yields
$L\!\left(S(3z+\tau;3\tau)\right)=-\sqrt{6}v^{\frac{3}{2}}q^{\frac{1}{6}}\zeta e^{-\frac{6\pi y^{2}}{v}}\sum_{n\in\mathbb{Z}-\frac{1}{6}}\left(n-\frac{y}{v}\right)(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}}\overline{\zeta}^{3n}.$
Next, we turn to the minus sign. Using that $S$ is an even function and changing $z\mapsto-z$ in the above, yields
$L\!\left(S(3z-\tau;3\tau)\right)=L\!\left(S(-3z+\tau;3\tau)\right)$ $=-\sqrt{6}v^{\frac{3}{2}}q^{\frac{1}{6}}\zeta^{-1}e^{-\frac{6\pi y^{2}}{v}}\!\!\!\sum_{n\in\mathbb{Z}-\frac{1}{6}}\!\!\left(n+\frac{y}{v}\right)(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}}\overline{\zeta}^{-3n}.$
Plugging the above two equations back into (3.12) gives
$L\!\left(R^{*}(z;\tau)\right)$ $=-\sqrt{\frac{3}{2}}v^{\frac{3}{2}}e^{-\frac{6\pi y^{2}}{v}}q^{\frac{1}{24}}\left(\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}\right)\sum_{\pm}\pm\!\sum_{n\in\mathbb{Z}-\frac{1}{6}}\left(n\mp\frac{y}{v}\right)(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}}\overline{\zeta}^{\pm 3n}$
$=-i\sqrt{6}v^{\frac{3}{2}}e^{\frac{3\pi(z-\overline{z})^{2}}{2v}}q^{\frac{1}{24}}\sin(\pi z)\sum_{\pm}\pm\sum_{n\in\mathbb{Z}-\frac{1}{6}}\left(n\mp\frac{z-\overline{z}}{2iv}\right)(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}}\overline{\zeta}^{\pm 3n},$
plugging in $y=\frac{z-\overline{z}}{2i}$. Using (3.11) and (3.6), we obtain
$\sum_{k\geq 0}\left(2L\!\left(\widehat{f}_{k}(\tau)\right)+\frac{3}{4\pi}\delta_{k=2}\right)\!\frac{(2\pi iz)^{k}}{k!}=\frac{L\!\left(R^{*}(z;\tau)\right)}{R^{*}(z;\tau)}\\ =-\frac{i\sqrt{6}v^{\frac{3}{2}}e^{\frac{3\pi(z-\overline{z})^{2}}{2v}}q^{\frac{1}{24}}\sin(\pi z)}{R^{*}(z;\tau)}\sum_{\pm}\pm\sum_{n\in\mathbb{Z}-\frac{1}{6}}\left(n\mp\frac{z-\overline{z}}{2iv}\right)(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}}\overline{\zeta}^{\pm 3n}.$
Collecting the constant terms in the Taylor expansions with respect to $\overline{z}$ on both sides yields
$\sum_{k\geq 0}\!\left(\!2L\Bigl{(}\widehat{f}_{k}(\tau)\Bigr{)}\!+\!\frac{3}{4\pi}\delta_{k=2}\!\right)\!\frac{(2\pi iz)^{k}}{k!}\!=\!-\frac{i\sqrt{6}v^{\frac{3}{2}}e^{\frac{3\pi z^{2}}{2v}}q^{\frac{1}{24}}\sin(\pi z)}{R^{*}(z;\tau)}\sum_{\pm}\pm\!\!\!\sum_{n\in\mathbb{Z}-\frac{1}{6}}\!\!\!\left(n\!\mp\!\frac{z}{2iv}\right)\!(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}}$
$\hskip 56.9055pt=\frac{\sqrt{6v}ze^{\frac{3\pi z^{2}}{2v}}q^{\frac{1}{24}}\sin(\pi z)}{R^{*}(z;\tau)}\!\sum_{n\in\mathbb{Z}-\frac{1}{6}}\!(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}}=-iz\sqrt{\frac{3v}{2}}\frac{\eta(\tau)}{\widehat{\mathbb{F}}(z;\tau)}\!\sum_{n\in\mathbb{Z}-\frac{1}{6}}\!(-1)^{n-\frac{5}{6}}\overline{q}^{\frac{3n^{2}}{2}},$
using (3.6). The sum on the right-hand side equals $-\eta(-\overline{\tau})$, so the above becomes
$iz\sqrt{\frac{3v}{2}}\frac{|\eta(\tau)|^{2}}{\widehat{\mathbb{F}}(z;\tau)}=-\pi\sqrt{6v}|\eta(\tau)|^{2}z^{2}\exp\!\left(-2\sum_{k\geq 1}\widehat{f}_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right),$
by (3.6) again. Using Lemma 2.7 with $w=2\pi iz$ and $x_{k}=-\frac{2}{k!}\widehat{f}_{k}(\tau)$, we conclude
$\sum_{k\geq 0}\!\left(\!2L\Bigl{(}\widehat{f}_{k}(\tau)\Bigr{)}+\frac{3}{4\pi}\delta_{k=2}\!\right)\!\frac{(2\pi iz)^{k}}{k!}=\frac{\sqrt{3v}}{2\sqrt{2}\pi}|\eta(\tau)|^{2}(2\pi iz)^{2}\sum_{n\geq 0}\mathrm{Tr}_{n}\bigl{(}\psi,\!\widehat{f};\!\tau\bigr{)}(2\pi iz)^{n}\!.$
Comparing the coefficient of $(2\pi iz)^{k}$ gives the claim for $k\geq 2$.
(4) The Serre derivative $\vartheta_{k}$ defined in (2.4) increases the weight of a modular object by $2$ (note that the proof in [17, p. 48] goes through if $f$ is non-holomorphic). We have
$D+\frac{2k}{3}f_{2}^{*}-\vartheta_{k}=2k\left(\frac{1}{3}f_{2}^{*}-G_{2}\right).$
By Lemma 3.3 and (1.4) this is a non-holomorphic modular form of weight $2$. Hence $D+\frac{2k}{3}f_{2}^{*}$ maps a modular object of weight $k$ to a modular object of weight $k+2$. ∎
We are now ready to prove Theorem 1.2.
Proof of Theorem 1.2.
Using (3.1) and Lemma 2.7 with $w=2\pi iz$ and $x_{k}=\frac{2}{k!}f_{k}$, we find
$R(\zeta;q)=\frac{\sin\!\left(\pi z\right)}{\pi z(q)_{\infty}}\sum_{k\geq 0}\mathrm{Tr}_{k}(\phi,f;\tau)(2\pi iz)^{k}.$
Expanding $R(\zeta;q)=\sum_{k\geq 0}R_{k}(q)\frac{(2\pi iz)^{k}}{k!}$, letting $z\mapsto\frac{z}{2\pi i}$, and using that $\sin(-\frac{iz}{2})=-i\sinh(\frac{z}{2})$, we conclude
$\sum_{k\geq 0}R_{k}(q)\frac{z^{k}}{k!}=\frac{2\sinh\!\left(\frac{z}{2}\right)}{z(q)_{\infty}}\sum_{k\geq 0}\mathrm{Tr}_{k}(\phi,f;\tau)z^{k}.$
By Lemma 3.5 the first property is satisfied, by Lemma 3.3 the second, and by Theorem 1.6 (1) the third. ∎
4. Proof of Theorems 1.4 and 1.5
The following lemma rewrites the rank moments $R_{k}$ in terms of the $g_{\ell}$, defined by (1.7).
Lemma 4.1.
For $k\geq 1$, we have
[…]
Proof.
Both sides of the lemma are 00 for $k$ odd. Namely, since $N(m,n)=N(-m,n)$, we have that $R_{k}$ is 00 for $k$ odd, and for $\ell$ odd, we have
$\left(2^{\ell-1}-1\right)\frac{B_{\ell}}{2\ell}+g_{\ell}(\tau)=0.$ (4.1)
So, we may assume that $k$ is even. Recall that by [5, equation (2.12)] we have, for $k\geq 2$ even,
$R_{k}(q)=\frac{2}{(q)_{\infty}}\sum_{n\geq 1}(-1)^{n+1}q^{\frac{n(3n-1)}{2}}(1-q^{n})\sum_{m\geq 0}m^{k}q^{nm}.$ (4.2)
Distinguishing between $n$ even and $n$ odd yields
$2^{k-1}(q)_{\infty}R_{k}(q)$ $\!=\!\!\!\!\sum_{n\geq 1,m\geq 0}\!\!\!(2m)^{k}\!\left(\!-q^{n(6n+2m-1)}\!+\!q^{(2n-1)(3n+m-2)}\!+\!q^{n(6n+2m+1)}\!-\!q^{(2n-1)(3n+m-1)}\!\right)$
$\!=\!\sum_{m\geq 3n\geq 3}(2m-6n)^{k}\!\left(\!-q^{n(2m-1)}+q^{(2n-1)(m-2)}+q^{n(2m+1)}-q^{(2n-1)(m-1)}\!\right)\!,$
making the change of variables $m\mapsto m-3n$. Interchanging the role of $m$ and $n$ in the second and fourth sum the above becomes
$-\sum_{m\geq 3n\geq 3}(2m-6n)^{k}q^{n(2m-1)}+\sum_{n\geq 3m\geq 3}(2n-6m)^{k}q^{(2m-1)(n-2)}\\ +\sum_{m\geq 3n\geq 3}(2m-6n)^{k}q^{n(2m+1)}-\sum_{n\geq 3m\geq 3}(2n-6m)^{k}q^{(2m-1)(n-1)}.$
Making the change of variables $n\mapsto n+2$ in the second sum, the change of variables $m\mapsto m-1$ in the third sum, and the change of variables $n\mapsto n+1$ in the fourth sum, the above becomes
$-\sum_{2m-1\geq 6n-1\geq 5}(2m-1-6n+1)^{k}q^{n(2m-1)}+\sum_{2n+1\geq 3(2m-1)\geq 3}(2n+1-3(2m-1))^{k}q^{n(2m-1)}$
$+\sum_{2m-1\geq 6n+1\geq 7}(2m-1-6n-1)^{k}q^{n(2m-1)}-\sum_{2n-1\geq 3(2m-1)\geq 3}(2n-1-3(2m-1))^{k}q^{n(2m-1)}$
$\qquad=\sum_{\begin{subarray}{c}m\geq 6n+1\geq 7\\ \text{$m$ odd}\end{subarray}}\left((m-6n-1)^{k}-(m-6n+1)^{k}\right)q^{nm}$
$\hskip 142.26378pt+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\\ \text{$m$ odd}\end{subarray}}\left((2n-3m+1)^{k}-(2n-3m-1)^{k}\right)q^{nm}$
$\qquad=\sum_{\begin{subarray}{c}m\geq 6n+1\geq 7\end{subarray}}\left((m-6n-1)^{k}-(m-6n+1)^{k}\right)q^{nm}$
$\hskip 142.26378pt+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\end{subarray}}\left((2n-3m+1)^{k}-(2n-3m-1)^{k}\right)q^{nm}.$ (4.3)
We require the identity
$(x+1)^{k}-(x-1)^{k}=2\sum_{\begin{subarray}{c}1\leq\ell\leq k-1\\ 2\nmid\ell\end{subarray}}\binom{k}{\ell}x^{\ell}=2\sum_{\begin{subarray}{c}2\leq\ell\leq k\\ 2\mid\ell\end{subarray}}\binom{k}{\ell-1}x^{\ell-1}.$
Using this and the fact that $k$ is even, we obtain that (4.3) equals
$2\sum_{\begin{subarray}{c}\ell=2\\ 2\mid\ell\end{subarray}}^{k}\binom{k}{\ell-1}\left(-\sum_{\begin{subarray}{c}m\geq 6n+1\geq 7\end{subarray}}(m-6n)^{\ell-1}q^{nm}+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\end{subarray}}(2n-3m)^{\ell-1}q^{nm}\right).$
Now the lemma follows from the definition of $g_{\ell}$. ∎
The generating function of the $g_{\ell}$ is closely related to the rank generating function.
Lemma 4.2.
We have
$\frac{\pi z(q)_{\infty}}{\sin\left(\pi z\right)}R(\zeta;q)\\ =1+4\sum_{k\geq 1}g_{k}(\tau)\frac{(\pi iz)^{k}}{(k-1)!}.$
Proof.
From (1.1) and (1.5) we deduce that
$R(\zeta;q)=\sum_{k\geq 0}R_{k}(q)\frac{(2\pi iz)^{k}}{k!}.$ (4.4)
By setting $X=\frac{1}{2}$ and $t=-2\pi iz$ in (2.5) we have
$\frac{\pi z}{\sin(\pi z)}$ $=\sum_{n\geq 0}B_{n}\!\left(\frac{1}{2}\right)\frac{(2\pi iz)^{n}}{n!}.$ (4.5)
Note that, by (1.5), $R_{0}(q)=\sum_{n\geq 0}\sum_{m\in\mathbb{Z}}N(m,n)q^{n}=\frac{1}{(q)_{\infty}}$ is the generating function of partitions. Hence, using (4.4) and Lemma 4.1, we obtain
$\frac{\pi z(q)_{\infty}}{\sin\left(\pi z\right)}\left(R(\zeta;q)-\frac{1}{(q)_{\infty}}\right)$
[…]
where we make the change of variables $n\mapsto n-k$. Using (2.6), (2.7), and the fact that $B_{m}(\frac{1}{2})=0$ and $B_{m}(0)=-\frac{\delta_{m=1}}{2}$ for $m$ odd, for $\ell\in\mathbb{N}$ odd and $n\in\mathbb{N}$ we find
$\sum_{\begin{subarray}{c}k=\ell\\ k\not\equiv\ell\ \,(\mathrm{mod}\,2)\end{subarray}}^{n}{n\choose k}\binom{k}{\ell}B_{n-k}\!\left(\frac{1}{2}\right)2^{\ell-k}=-\binom{n}{\ell}B_{n-\ell}(0)=\begin{cases}\frac{n}{2}&\text{if }n=\ell+1,\\ 0&\text{otherwise}.\end{cases}$
Using the above and (4.1), we find
$\frac{\pi z(q)_{\infty}}{\sin\left(\pi z\right)}\left(R(\zeta;q)-\frac{1}{(q)_{\infty}}\right)=\sum_{n\geq 1}\frac{n}{2^{n-2}}\left(\frac{\left(2^{n-1}-1\right)B_{n}}{2n}+g_{n}(\tau)\right)\frac{(2\pi iz)^{n}}{n!}.$
The result follows by adding $\frac{\pi z}{\sin(\pi z)}$ on both sides and using (4.5) and the fact that $B_{n}(\frac{1}{2})=-(1-2^{1-n})B_{n}$. ∎
Now, we are ready to prove Theorems 1.4 and 1.5.
Proof of Theorem 1.4.
(1) Using Lemma 4.2 and then taking the derivative of (3.1) with respect to $z$ yields
$4\pi i\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k-1}}{(k-1)!}\exp\!\left(\!2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\!\right)=2\pi i\sum_{k\geq 0}\frac{kg_{k}(\tau)}{2^{k-2}}\frac{(2\pi iz)^{k-1}}{(k-1)!}.$ (4.6)
Again using Lemma 4.2 we obtain
$2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k-1}}{(k-1)!}+2\sum_{n,\ell\geq 1}\binom{n+\ell-1}{n-1}f_{n}(\tau)\frac{\ell g_{\ell}(\tau)}{2^{\ell-2}}\frac{(2\pi iz)^{n+\ell-1}}{(n+\ell-1)!}=\sum_{k\geq 0}\frac{kg_{k}(\tau)}{2^{k-2}}\frac{(2\pi iz)^{k-1}}{(k-1)!}.$
By extracting the coefficients of $(2\pi iz)^{k-1}$, we obtain (1).
(2) We send the exponential in (4.6) to the other side and apply Lemma 2.7 with $w=2\pi iz$ and $x_{k}=-\frac{2f_{k}}{k!}$ to obtain
$2\sum_{k\geq 1}f_{k}(\tau)\frac{(2\pi iz)^{k-1}}{(k-1)!}=\sum_{n\geq 0}\frac{ng_{n}(\tau)}{2^{n-2}}\frac{(2\pi iz)^{n-1}}{(n-1)!}\sum_{m\geq 0}\mathrm{Tr}_{m}(\psi,f;\tau){(2\pi iz)^{m}}.$
Comparing the coefficients of $(2\pi iz)^{k-1}$ gives (2). ∎
Proof of Theorem 1.5.
By (3.1), (4.4), and Lemma 2.3, we find that
$\exp\!\left(2\sum_{k\geq 1}\left(f_{k}(\tau)+\frac{B_{k}}{2k}\right)\frac{z^{k}}{k!}\right)$ $=1+(q)_{\infty}\sum_{k\geq 1}R_{k}(q)\frac{z^{k}}{k!}.$
Taking the logarithm and expanding the right-hand side formally, we deduce that
$2\sum_{k\geq 1}\left(f_{k}(\tau)+\frac{B_{k}}{2k}\right)\frac{z^{k}}{k!}=\sum_{n\geq 1}\frac{(-1)^{n+1}}{n}\sum_{k_{1},\ldots,k_{n}\geq 1}z^{k_{1}+\ldots+k_{n}}\prod_{j=1}^{n}(q)_{\infty}R_{k_{j}}(q)\frac{1}{k_{j}!}.$
Let $\Omega$ be the linear map on $\mathbb{C}[\![z]\!]$ given by $\Omega(z^{k}):=k!z^{k}$ for $k\geq 0$. Applying $\Omega$, we obtain
$\!\!2\sum_{k\geq 1}\!\left(f_{k}(\tau)\!+\!\frac{B_{k}}{2k}\right)\!z^{k}$ $=\sum_{n\geq 1}\frac{(-1)^{n+1}}{n}\mathcal{R}_{n}(z;q),$ (4.7)
where we set
$\mathcal{R}_{n}(z;q):=\sum_{k_{1},\ldots,k_{n}\geq 1}\!\!\binom{k_{1}+\ldots+k_{n}}{k_{1},\ldots,k_{n}}z^{k_{1}+\ldots+k_{n}}\prod_{j=1}^{n}(q)_{\infty}R_{k_{j}}(q).$
To show that the Fourier coefficients of $f_{k}(\tau)\!+\!\frac{B_{k}}{2k}$ are integral, it suffices to show that these on the right-hand side of (4.7) are even integers. By (4.2) the Fourier coefficients of $(q)_{\infty}R_{k}(q)$ are even. Hence, those of $\frac{1}{n}\mathcal{R}_{n}(z;q)$ are even. Let $k_{1},\ldots,k_{n}\in\mathbb{N}$. There exists a unique partition $\lambda$ of length $n$ associated to these $k_{j}$ (obtained by ordering the $k_{j}$ in non-increasing order). Conversely, for a partition $\lambda$ of length $n$, there are $\binom{n}{r_{1}(\lambda),r_{2}(\lambda),\ldots}$ many $n$-tuples $(k_{1},\ldots,k_{n})$ with associated partition $\lambda$, where $r_{m}(\lambda)$ denotes the number of parts of size $m$ in $\lambda$. Let $\mathcal{P}$ be the set of partitions and $|\lambda|=\sum_{j}\lambda_{j}$ the size of $\lambda\in\mathcal{P}$. We replace the sum over $k_{1},\ldots,k_{n}$ by a sum over partitions. Then, we have $\binom{k_{1}+\ldots+k_{n}}{k_{1},\ldots,k_{n}}=\binom{|\lambda|}{\lambda_{1},\ldots,\lambda_{n}}$ and
$\mathcal{R}_{n}(z;q)=\sum_{\begin{subarray}{c}\lambda\in\mathcal{P}\\ \ell(\lambda)=n\end{subarray}}{\binom{|\lambda|}{\lambda_{1},\ldots,\lambda_{n}}}\binom{n}{r_{1}(\lambda),r_{2}(\lambda),\ldots}z^{|\lambda|}\prod_{j=1}^{n}(q)_{\infty}R_{\lambda_{j}}(q).$
The claim follows once we show that
$n\,\Big{|}\,\binom{|\lambda|}{\lambda_{1},\ldots,\lambda_{n}}\binom{n}{r_{1}(\lambda),r_{2}(\lambda),\ldots}.$ (4.8)
Let $d$ be the greatest common divisor of $r_{1}(\lambda)$, $r_{2}(\lambda),\ldots$. Since $\sum_{m\geq 1}r_{m}(\lambda)=n$, we have that $d\mid n$ and we write $n=rd$. Moreover, note that one can write the partition $(\lambda_{1},\ldots,\lambda_{n})$ as $(\ell_{1},\ldots,\ell_{1},\ell_{2},\ldots,\ell_{2},\ldots,\ell_{r},\ldots,\ell_{r})$, where each $\ell_{j}$ is repeated $d$ times. First, we show that
$d\,\Big{|}\,\binom{|\lambda|}{\lambda_{1},\ldots,\lambda_{n}}.$ (4.9)
We factorize the multinomial coefficient as
$\binom{|\lambda|}{\lambda_{1},\ldots,\lambda_{n}}=\binom{|\lambda|}{\tfrac{|\lambda|}{d},\ldots,\tfrac{|\lambda|}{d}}\binom{\frac{|\lambda|}{d}}{\ell_{1},\ldots,\ell_{r}}^{\!d},$ (4.10)
where $\tfrac{|\lambda|}{d}$ occurs $d$ times in the first multinomial coefficient on the right-hand side. All of the multinomial coefficients are integers and by Lemma 2.8 with $n=|\lambda|$ and $a_{j}=\frac{|\lambda|}{d}$ for $j\in\{1,\ldots,d\}$, we find that $d$ divides the first factor in (4.10) and hence, (4.9) holds.
Again using Lemma 2.8, in this case with $a_{m}=r_{m}(\lambda)$ and $d=\gcd(a_{1},\ldots,a_{\ell})$, we have $r\,|\,\binom{n}{r_{1}(\lambda),r_{2}(\lambda),\ldots}.$ Combining this with (4.9) gives (4.8). ∎
Finally, we refine Lemma 3.5 by computing the first Fourier coefficients of $f_{k}$. In particular, the $f_{k}$ are naturally normalized in the sense that the second Fourier coefficient is $1$.
Proposition 4.3.
For $k\geq 2$ even we have
$f_{k}(\tau)=-\frac{B_{k}}{2k}+q^{2}+\left(2^{k}-1\right)q^{3}+O\left(q^{4}\right).$
Proof.
Recall that $(q)_{\infty}R_{k}(q)=0$ for $k$ odd. By (4.2), we have for $k$ even
$(q)_{\infty}R_{k}(q)=2q^{2}+2\left(2^{k}-1\right)q^{3}+O\left(q^{4}\right).$
Hence, we have $\prod_{j=1}^{n}(q)_{\infty}R_{k_{j}}(q)=O(q^{4})$ if $n\geq 2$ and $k_{j}\in\mathbb{N}$ even. Therefore, by (4.7),
$2\sum_{k\geq 1}\left(f_{k}(\tau)+\frac{B_{k}}{2k}\right){z^{k}}$ $=2\sum_{\begin{subarray}{c}k\geq 2\\ k\text{ even}\end{subarray}}\!\left(q^{2}+\left(2^{k}-1\right)q^{3}\right)z^{k}+O\!\left(q^{4}\right).\qed$
5. Examples
Here, we write down the first Fourier coefficients of the mock Eisenstein series $f_{k}$. We have
$f_{2}(\tau)$ $=-\frac{1}{24}+q^{2}+3q^{3}+5q^{4}+7q^{5}+9q^{6}+10q^{7}+13q^{8}+O\!\left(q^{9}\right)\!,$
$f_{4}(\tau)$ $=\frac{1}{240}+q^{2}+15q^{3}+59q^{4}+139q^{5}+255q^{6}+406q^{7}+595q^{8}+O\!\left(q^{9}\right)\!,$
$f_{6}(\tau)$ $=-\frac{1}{504}+q^{2}+63q^{3}+635q^{4}+2827q^{5}+8199q^{6}+18550q^{7}+36043q^{8}+O\!\left(q^{9}\right)\!,$
$f_{8}(\tau)$ $=\frac{1}{480}+q^{2}+255q^{3}+6179q^{4}+53179q^{5}+253815q^{6}+844966q^{7}+2234875q^{8}+O\!\left(q^{9}\right)\!.$
Some examples for Theorem 1.6 (1) are given by
$D(f_{2})$ $=-f_{2}G_{2}-\frac{f_{2}^{2}}{2}-\frac{f_{4}}{12}+\frac{3G_{2}^{2}}{2}+\frac{G_{4}}{12},$
$D(f_{4})$ $=6f_{2}^{2}G_{2}-18f_{2}G_{2}^{2}-f_{2}G_{4}-f_{4}G_{2}-\frac{2}{3}f_{2}^{3}-\frac{7f_{4}f_{2}}{3}-\frac{f_{6}}{9}+18G_{2}^{3}+3G_{2}G_{4}+\frac{G_{6}}{30},$
$D(f_{6})$ $=-60f_{2}^{3}G_{2}+270f_{2}^{2}G_{2}^{2}+15f_{2}^{2}G_{4}-540f_{2}G_{2}^{3}+30f_{4}f_{2}G_{2}-90f_{2}G_{2}G_{4}-f_{2}G_{6}$
$\hskip 14.22636pt-45f_{4}G_{2}^{2}-f_{6}G_{2}-\frac{5f_{4}G_{4}}{2}+5f_{2}^{4}-5f_{4}f_{2}^{2}-\frac{11f_{6}f_{2}}{3}-\frac{25f_{4}^{2}}{4}-\frac{f_{8}}{8}+405G_{2}^{4}$
$\hskip 14.22636pt+\frac{21855G_{4}^{2}}{3652}+135G_{2}^{2}G_{4}+3G_{2}G_{6}-\frac{39G_{8}}{51128},$
$D(f_{8})$ $=840f_{2}^{4}G_{2}-5040f_{2}^{3}G_{2}^{2}-280f_{2}^{3}G_{4}+15120f_{2}^{2}G_{2}^{3}-840f_{4}f_{2}^{2}G_{2}+2520f_{2}^{2}G_{2}G_{4}$
$\hskip 14.22636pt+28f_{2}^{2}G_{6}-22680f_{2}G_{2}^{4}+2520f_{4}f_{2}G_{2}^{2}-\frac{305970}{913}f_{2}G_{4}^{2}+56f_{6}f_{2}G_{2}-7560f_{2}G_{2}^{2}G_{4}$
$\hskip 14.22636pt+140f_{4}f_{2}G_{4}-168f_{2}G_{2}G_{6}+\frac{39f_{2}G_{8}}{913}-2520f_{4}G_{2}^{3}-84f_{6}G_{2}^{2}+70f_{4}^{2}G_{2}-f_{8}G_{2}$
$\hskip 14.22636pt-420f_{4}G_{2}G_{4}-\frac{14f_{6}G_{4}}{3}-\frac{14f_{4}G_{6}}{3}-56f_{2}^{5}+\frac{280}{3}f_{4}f_{2}^{3}-\frac{28}{3}f_{6}f_{2}^{2}-\frac{70}{3}f_{4}^{2}f_{2}-5f_{8}f_{2}$
$\hskip 14.22636pt-\frac{322f_{4}f_{6}}{9}-\frac{2f_{10}}{15}+13608G_{2}^{5}+\frac{917910}{913}G_{2}G_{4}^{2}+7560G_{2}^{3}G_{4}+252G_{2}^{2}G_{6}$
$\hskip 14.22636pt+\frac{19352886G_{4}G_{6}}{1983949}+\frac{36751G_{10}}{1803590}-\frac{117G_{2}G_{8}}{913}.$
6. Questions for future research
We end by raising some open questions.
  1. (1)
    The three properties of the $f_{k}$ given in Theorem 1.2 do not determine them uniquely. Therefore, we provide two recursive definitions of $f_{k}$ in Theorem 1.4. It would be interesting to find “nice” properties that define a mock Eisenstein series uniquely. For example, the fact that $G_{k}$ is a normalized Hecke eigenform of weight $k$ which does not vanish at $i\infty$ determines it uniquely. As the coefficient of $q$ in the Fourier expansion of the $f_{k}$ vanishes by Proposition 4.3, one deduces that $f_{k}$ cannot be a Hecke eigenform.
  2. (2)
    The functions $f_{2},f_{4},f_{6},\ldots$ together with $G_{2},G_{4}$ and $G_{6}$, do not seem to satisfy any algebraic relations. This has been verified numerically up to weight 24 (and also up to mixed weight 12). Is it indeed the case that the algebra $\mathcal{F}$ is free?
  3. (3)
    What variations of the $g_{\ell}$, in particular in the range of summation, are also mock modular? More concretely, for which $a$, $b\in\mathbb{N}$ is the function
    $\sum_{\begin{subarray}{c}an-1\geq bm\geq b\end{subarray}}(an-bm)^{\ell-1}q^{mn}-\sum_{\begin{subarray}{c}n-1\geq abm\geq ab\end{subarray}}(n-abm)^{\ell-1}q^{mn}$
    a mock Eisenstein series? Note that $a=2$ and $b=3$ yields the function $g_{\ell}$ in this paper. Are other choices of $a$ and $b$ also of particular interest? Do these functions, together with those in [18, p.15] and [13, equation (1.4)], form the first examples of a theory of (higher level) mock Eisenstein series?